# Two-Point Stabilizer Rényi Entropy: a Computable Magic Proxy of Interacting Fermions

### v1

@@ Abstract
$M_{\alpha}(\rho)=\frac{1}{1-\alpha}\log\sum_{P\in\mathcal{P}_{N}}\frac{1}{d}\left[\text{Tr}(P\rho)\right]^{\alpha}-S_{\alpha}(\rho),$ (1)
@@
We revisit the one-dimensional half-filled $t$-$V$ model [62, 63, 64], utilizing DMRG simulations [65] with open boundary conditions (centering the reference site $i$). We first analyze the spatial scaling behavior as the distance |ij||i-j| varies. According to bosonization theory, the single particle Green’s function $\langle c_{i}^{\dagger}c_{i+r}\rangle$ decays algebraically as $r^{-\eta}$ (with $\eta\approx 1$) in the Luttinger liquid (LL) phase ($V<2$), whereas it exhibits exponential decay $e^{-r/\xi}$ in the gapped charge density wave (CDW) phase. Leveraging tomography techniques [66], we derive the general analytical form of the fermionic two-point SRE. To a first order approximation, this quantity is dominated by the squared magnitude of the Green’s function, superimposed onto a background of intrinsic magic density. In the LL regime, the single site reduced density matrix corresponds to a mixed stabilizer state, yielding zero intrinsic magic. In contrast, the CDW phase retains a finite magic density arising from spontaneous symmetry breaking. To isolate the correlation dependent scaling from this local background, we introduce the mutual two-point SRE, defined as $\tilde{\mathcal{M}}^{(\alpha)}_{i,j}(\rho)=\mathcal{M}_{i,j}^{(\alpha)}(\rho)-\mathcal{M}_{i}^{(\alpha)}(\rho)-\mathcal{M}_{j}^{(\alpha)}(\rho)$. Figure 2(a) displays the computed mutual two-point SRE across distinct phases. The numerical results corroborate our analytical predictions. Specifically, in the interacting LL phase, we observe that the power law decay exponent deviates slightly from the non-interacting limit of $2$, reflecting interaction effects. Subsequently, we confirm that this exponent modification, and the correlation length $\xi(V)$, can be precisely extracted via the susceptibility $\partial\tilde{\mathcal{M}}/\partial V$, providing an effective scheme for characterizing the phase transition.
@@
Upon tuning the interaction strength $V$, $\tilde{\mathcal{M}}^{(2)}_{i,j}$ emerges as a sensitive indicator of the quantum phase transition between the LL and CDW phases. As illustrated in Fig. 2(b), the mutual two-point SRE displays a pronounced peak characterized by a discontinuity in its first derivative at the critical point, a feature that becomes sharper as the system size increases. We attribute this behavior to the distinct scaling laws governing $\tilde{\mathcal{M}}$ in the respective phases. In the gapless LL regime, the algebraic decay yields a derivative scaling of $\frac{\partial\mathcal{M}_{i,i+r}^{(2)}}{\partial V}\sim-2\frac{\partial\alpha}{\partial V}r^{-2\alpha(V)}\ln r$. Conversely, in the gapped CDW phase, the exponential decay leads to $\frac{\partial\mathcal{M}_{i,i+r}^{(2)}}{\partial V}\sim\frac{r}{\xi^{2}(V)}e^{-r/\xi(V)}\frac{\partial\xi}{\partial V}$. We identify the location of this maximum as the finite-size critical point, $V_{c}(L)$. Notably, the shift of $V_{c}(L)$ from the asymptotic value $V_{c}(\infty)=2$ follows the characteristic finite size scaling of a Berezinskii-Kosterlitz-Thouless (BKT) transition [67, 68, 69, 70, 71, 72], where the correlation length diverges as $\xi\sim\exp{\left(\pi^{2}/2\sqrt{|V-V_{c}|}\right)}$. Equating the correlation length to the system size, $\xi\sim L$, yields the scaling relation $V_{c}(L)=2+\frac{\pi^{4}/4}{(\ln L-\ln L_{0})^{2}}$ where $L_{0}$ is a non-universal microscopic length scale. As shown in Fig. 2(c), $V_{c}(L)$ is in excellent agreement with this scaling relation. Therefore, we conclude that the two-point SRE accurately captures the critical scaling properties of the transition.
@@
To circumvent the uncertainty in $V_{c}$, we perform the scaling analysis against the dimensionless correlation ratio $R=1-\frac{C(k=q_{\text{min}})}{C(k=0)}$ [76], where $C(k)$ is the Fourier transform of the density-density correlation and $q_{\text{min}}=(2\pi/L,0)$. Using the package in Ref. [77], we determine $\eta\approx 0.423$, which is consistent with Gross-Neveu-Ising theory [78], achieving optimal data collapse for the scaled $\tilde{\mathcal{M}}^{(2)}_{i,i+\vec{r}_{\max}}$ versus $R$, as shown in Fig. 3(b). This result further confirms the capability of the mutual two-point SRE to accurately determine critical exponents at two-dimensional quantum critical points.
@@
1D $t$-$V$ Model: In the Luttinger liquid (LL) phase at half-filling ($n_{i}=0.5,m_{i}=0$), the mutual SRE is dominated by the Green’s function, scaling as $\tilde{\mathcal{M}}^{(2)}\sim|g_{r}|^{2}\sim r^{-2\eta}$ (where $|g_{r}|\sim r^{-\eta}$). In the CDW phase, the energy gap dictates exponential decay, $\tilde{\mathcal{M}}^{(2)}\sim e^{-r/\xi}$. The distinct scaling of the susceptibility $\partial_{V}\tilde{\mathcal{M}}^{(2)}$—diverging as $2\partial_{V}\alpha r^{-2\alpha}\ln r$ in the LL phase versus peaking as $\xi^{-2}re^{-r/\xi}\partial_{V}\xi$ in the CDW phase—signals the phase transition. Finite-size effects shift the peak of the latter to $V_{c}(L)$, governed by the BKT-like correlation length scaling $\xi\sim\exp(\pi^{2}/2\sqrt{|V-V_{c}|})$, leading to $V_{c}(L)\approx 2+\pi^{4}/4(\ln L-\ln L_{0})^{2}$.

### v2

@@ Abstract
$M_{\alpha}(\rho)=\frac{1}{1-\alpha}\log\sum_{P\in\mathcal{P}_{N}}\frac{1}{d}\left[\text{Tr}(P\rho)\right]^{2\alpha}-S_{\alpha}(\rho),$ (1)
@@
We revisit the one-dimensional half-filled $t$-$V$ model [62, 63, 64], utilizing DMRG simulations [65] with open boundary conditions (centering the reference site $i$). We first analyze the spatial scaling behavior as the distance |ij||i-j| varies. According to bosonization theory, the single particle Green’s function $\langle c_{i}^{\dagger}c_{i+r}\rangle$ decays algebraically as $r^{-\eta}$ (with $\eta\approx 1$) in the Luttinger liquid (LL) phase ($V<2$), whereas it exhibits exponential decay $e^{-r/\xi}$ in the gapped charge density wave (CDW) phase. Leveraging tomography techniques [66], we derive the general analytical form of the fermionic two-point SRE. To a first order approximation, this quantity is dominated by the squared magnitude of the Green’s function, superimposed onto a background of intrinsic magic density. In the LL regime, the single site reduced density matrix corresponds to a mixed stabilizer state, yielding zero intrinsic magic. In contrast, the CDW phase retains a finite magic density arising from spontaneous symmetry breaking. To isolate the correlation dependent scaling from this local background, we introduce the mutual two-point SRE, defined as $\tilde{\mathcal{M}}^{(\alpha)}_{i,j}(\rho)=\mathcal{M}_{i,j}^{(\alpha)}(\rho)-\mathcal{M}_{i}^{(\alpha)}(\rho)-\mathcal{M}_{j}^{(\alpha)}(\rho)$. Figure 2(a) displays the computed mutual two-point SRE across distinct phases. The numerical results corroborate our analytical predictions. Specifically, in the interacting LL phase, we observe that the power law decay exponent deviates slightly from the non-interacting limit of $2$, reflecting interaction effects. Subsequently, we confirm that this exponent modification, and the correlation length $\xi$, can be precisely extracted via the susceptibility $\partial\tilde{\mathcal{M}}/\partial V$, providing an effective scheme for characterizing the phase transition.
@@
Upon tuning the interaction strength $V$, $\tilde{\mathcal{M}}^{(2)}_{i,j}$ emerges as a sensitive indicator of the quantum phase transition between the LL and CDW phases. As illustrated in Fig. 2(b), the mutual two-point SRE displays a pronounced peak characterized by a discontinuity in its first derivative at the critical point, a feature that becomes sharper as the system size increases. We attribute this behavior to the distinct scaling laws governing $\tilde{\mathcal{M}}$ in the respective phases. In the gapless LL regime, the algebraic decay yields a derivative scaling of $\frac{\partial\mathcal{M}_{i,i+r}^{(2)}}{\partial V}\sim-2\frac{\partial\eta}{\partial V}r^{-2\eta}\ln r$. Conversely, in the gapped CDW phase, the exponential decay leads to $\frac{\partial\mathcal{M}_{i,i+r}^{(2)}}{\partial V}\sim\frac{r}{\xi^{2}}e^{-r/\xi}\frac{\partial\xi}{\partial V}$. We identify the location of this maximum as the finite-size critical point, $V_{c}(L)$. Notably, the shift of $V_{c}(L)$ from the asymptotic value $V_{c}(\infty)=2$ follows the characteristic finite size scaling of a Berezinskii-Kosterlitz-Thouless (BKT) transition [67, 68, 69, 70, 71, 72], where the correlation length diverges as $\xi\sim\exp{\left(\pi^{2}/2\sqrt{|V-V_{c}|}\right)}$. Equating the correlation length to the system size, $\xi\sim L$, yields the scaling relation $V_{c}(L)=2+\frac{\pi^{4}/4}{(\ln L-\ln L_{0})^{2}}$ where $L_{0}$ is a non-universal microscopic length scale. As shown in Fig. 2(c), $V_{c}(L)$ is in excellent agreement with this scaling relation. Therefore, we conclude that the two-point SRE accurately captures the critical scaling properties of the transition.
@@
To circumvent the uncertainty in $V_{c}$, we perform the scaling analysis against the dimensionless correlation ratio $R=1-\frac{C(k=q_{\text{min}})}{C(k=0)}$ [76], where $C(k)$ is the Fourier transform of the density-density correlation and $q_{\text{min}}=(2\pi/L,0)$. Using the package in Ref. [77], we determine $\eta\approx 0.423$, which is consistent with Gross-Neveu-Ising theory [78], achieving optimal data collapse for the scaled $\tilde{\mathcal{M}}^{(2)}_{i,i+\vec{r}_{\max}}$ versus $R$, as shown in Fig. 3(b). Notably, analogous finite size scaling behavior has been reported for both density of magic and long range magic (a generalization of the mutual two point SRE) in spin systems [42, 79, 80, 45]. This result further confirms the capability of the mutual two-point SRE to accurately determine critical exponents at two-dimensional quantum critical points.
@@
1D $t$-$V$ Model: In the Luttinger liquid (LL) phase at half-filling ($n_{i}=0.5,m_{i}=0$), the mutual SRE is dominated by the Green’s function, scaling as $\tilde{\mathcal{M}}^{(2)}\sim|g_{r}|^{2}\sim r^{-2\eta}$ (where $|g_{r}|\sim r^{-\eta}$). In the CDW phase, the energy gap dictates exponential decay, $\tilde{\mathcal{M}}^{(2)}\sim e^{-r/\xi}$. The distinct scaling of the susceptibility $\partial_{V}\tilde{\mathcal{M}}^{(2)}$—diverging as $2(\partial_{V}\eta)r^{-2\eta}\ln r$ in the LL phase versus peaking as $\xi^{-2}re^{-r/\xi}\partial_{V}\xi$ in the CDW phase—signals the phase transition. Finite-size effects shift the peak of the latter to $V_{c}(L)$, governed by the BKT-like correlation length scaling $\xi\sim\exp(\pi^{2}/2\sqrt{|V-V_{c}|})$, leading to $V_{c}(L)\approx 2+\pi^{4}/4(\ln L-\ln L_{0})^{2}$.